A single cylinder single acting four stroke cycle gas engine develops 22 kw at 300 rpm the fly wheel mass is 100kg hoop stress developed is 5 Mpa density of material of rim of fly wheel is 800Kg/m 3.The speed variation on the either side is 1% of mean speed. Determine the ratio of work done during expansion and compression strokes work done in suction and exhaust stroke is negligible.

1
Expert's answer
2021-12-13T17:00:01-0500

"T_{mean}=\frac{P*6}{2π*N}"

"=\frac{22*10^3*60}{2*π*300}"

=700.28Nm

Work done per cycle="T_{mean}*\theta"

=700.28*4π

=2200

"Required\ ratio=\frac{W_e}{W_c}"

"=\frac{6418.38}{2381.62}"

=2.69